OpenStudy (anonymous):

Suppose that y = 21(1.07)x, represents book sales in billions of dollars since 2000. X represents the number of years after 2000. How many billions of dollars of books were sold in 2000? Use x = 0 and round the answer to the nearest whole number. Answer $21 billion $1.07 billion $22.47 billion $42 billion

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

Plug in x = 0 and evaluate to find y

OpenStudy (anonymous):

I got 22.47

jimthompson5910 (jim_thompson5910):

y = 21(1.07)^x y = 21(1.07)^0 y = ????

OpenStudy (anonymous):

21 billion

jimthompson5910 (jim_thompson5910):

correct